two investments totalling $55,000 produce an annual income of $2825. One investment yields 9% a year while the other yields 4% a year how much is invested at each rate? If i use the formula I=p.r.t I don't know what the principal or the rate is can you help me figure this out please?

Answer by checkley79(3341)

.09x+.04(55,000-x)=2,825
.09x+2,200-.04x=2,825
.05x=2,825-2,200
.05x=625
x=625/.05
x=$12,500 invested @ 9%.
55,000-12,500=$42,500 invested @ 4%.
Proof:
.09*12,500+.04*42,500=2,825
1,125+1,700=2,825
2,825=2,825